OpenStudy (anonymous):

I have some questions regarding combinations in finite. Here is my first one I am stumped on: Select 3 coins, but assume that there are 5 dimes, 3 nickels, and 2 quarters. In How many possible ways can the selection be made so that the value of the coins is AT LEAST 25 cents?_______________________________ Any help? Formulas have not done me so well

OpenStudy (anonymous):

i think you have to grind it til you find it

OpenStudy (anonymous):

sorry i was away

OpenStudy (anonymous):

oh good so i am less embarrassed. you have to work in cases

OpenStudy (anonymous):

i have tried a number of times but cant solve it

OpenStudy (anonymous):

cant break down the cases correctly

OpenStudy (anonymous):

i assume

OpenStudy (anonymous):

1) two quarters one dime only one way to select two quarters, 5 choices for the dime, so 5 in that case 2) two quarters one nickel, 3 choices for the nickel so 3 ways in that case

OpenStudy (anonymous):

3) one quarter two dimes \[2\times \dbinom{5}{2}=2\times 10=20\]

OpenStudy (anonymous):

4) one quarter two nickels \[2\times 3=6\] 5) one quarter one dime one nickel \[2\times 3\times 5=30\]

OpenStudy (anonymous):

and finally two dimes and a nickel \[10\times 3=30\] think that is all, so if we add them up we should get the right answer

OpenStudy (anonymous):

@zarkon, you got a snap way to do this, or it it just grind it out?

OpenStudy (anonymous):

wow i did not break the cases down enough I will check and see if it is right

OpenStudy (anonymous):

damn i forgot 3 dimes!!

OpenStudy (zarkon):

I do have a quicker way...but it is goog to do it this way...you also nee 3 dimes I believe

OpenStudy (anonymous):

I will try to add them up and see

OpenStudy (anonymous):

add another 10 what is the quick way for my benefit

OpenStudy (zarkon):

which gives a total of 104

OpenStudy (zarkon):

you can also look at it this way...

OpenStudy (anonymous):

now i get to learn something

OpenStudy (zarkon):

there are \[_{10}C_{3}=120\] ways to pick the 3 coins... only NNN and 2nickles and a dime are less than 25 cents 120-1-15=104

OpenStudy (anonymous):

yes much snappier

OpenStudy (zarkon):

I don't see a way to do it without some cases

OpenStudy (anonymous):

these few are killing me

OpenStudy (anonymous):

104 is correct

OpenStudy (anonymous):

yes well that would really be magic

OpenStudy (anonymous):

thank you so much I have some more that I am stumped as well if you want to crack them

OpenStudy (anonymous):

?:)

OpenStudy (anonymous):

post a new thread, you will get lots of answers i am sure